20210110 OSCO + INAF Full AO

01:05 OSCO Start
  • All Good
  • Take NCPA on and off data
  • looks elongated
01:55 Stop AO. Clouds coming in and Seeing go from ~1.1 to 2.1. Now back to 1.5 or so.
  • Tried bin 3 (accidently as Doug clicked Check Flux when he should not have)
  • Opitmize gain fail due to Skip Frame due to seeing from 1.1 => 2.1
  • WFS powered off
  • w_idle

-- DougMiller - 11 Jan 2021
Topic revision: r1 - 11 Jan 2021, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback